40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14617     Entry time: Fri May 17 10:57:01 2019
Author: gautam 
Type: Update 
Category: SUS 
Subject: IY chamber opened 

At ~930am, I vented the IY annulus by opening VAEV. I checked the particle count, seemed within the guidelines to allow door opening so I went ahead and loosened the bolts on the ITMY chamber.

Chub and I took the heavy door off with the vertex crane at ~1015am, and put the light door on.

Diagnosis plan is mainly inspection for now: take pictures of all OSEM/magnet positionings. Once we analyze those, we can decide which OSEMs we want to adjust in the holders (if any). I shut down the ITMY and SRM watchdogs in anticipation of in-chamber work.

Not related to this work: Since the annuli aren't being pumped on, the pressure has been slowly rising over the week. The unopened annuli are still at <1 torr, and the PAN region is at ~2 mtorr.

ELOG V3.1.3-